A simple pendulum is hanging from a support and is at rest as seen from the lab where the latitude is 600. Assume the earth to be a uniform sphere. The angle between the string and the true vertical at that place is

(a) less than 20


(b) between 20 and 40


(c) between 40 and 60


(d) more than 60?


1
Expert's answer
2021-03-29T09:04:23-0400

"T=m\omega^2 l=m \omega^2Rcos \varphi,"

"P=mg=m(\frac{GM}{R^2}-\omega^2 R),"

"F=\frac{GmM}{R^2},"

"\varphi=60°,"

"R=6.37\cdot 10^6~m,"

"\omega=\frac{2\pi}{t}=\frac{2\pi}{86164},"

"T^2=P^2+F^2-2PTcos \beta,\implies"

"cos\beta=\frac{P^2+F^2-T^2}{2PF}=\frac{(\frac{GM}{R^2}-\omega^2 R)^2+(\frac{GM}{R^2})^2-(\omega^2 R cos\varphi)^2}{2(\frac{GM}{R^2}-\omega^2 R)\frac{GM}{R^2}}\approx 1,"

"\beta\approx 0°<2°,"

answer (a).